1. Find the general solution to the next system of differential equations.

Y = [2 =3] [2 – 4) (1-3 y+ 2t2 + 10+] t2 +9t +3

2. Find the general solution of the following system of differential equations by parametric conversion.

Sa = - 3x+y+3t ly = 27 - 4y+et
Answer #1

1. ya 2 - 4 2+2 + 10+ 3) ý of - 3 +2 +95 +3 So y = 24, -4Y2 L1 + 2+2 +10t0 => (D-234, +4427 2t2 + lot l y = y, - 382 + +2 +9+ {ë at - t2 From (11) & y(t) = C et + Gelt, Therefore Y, Ye taip get 4Y₂ = 2t2 + 10t -Dy, +24, => 4 ₂ = 242 +10t. Get tale?+te + que C Geet & ost let) - èt 10 I 6-7 +10 whene , c2 ane constant let up be particular. solution. Then Alp (3+12+ + ēt) D2